Consider electrons accelerated using a potential difference of 25.5 kV before hitting a metal surface. Calculate the minimum wavelength of the emitted x-rays. Give your answer in pm and with 3 significant figures.
Answer #1

E = q* V
= (1.6*10^-19)*(25.5*10^3)
= 4.08*10^15 J

Now use:
E = h*c/wavelength
wavelength = h*c/E
= (6.626*10^-34)*(3*10^8)/(4.08*10^15)
= 4.87*10^-11 m
= 48.7 pm
Answer: 48.7 pm